OpenStudy (calculusxy):

Help on Word Problems! Krish has only 2 cups of diced pumpkins, when he needs 3. How many tablespoons of sugar must he use to keep the same proportions?

OpenStudy (calculusxy):

and sugar is 1 1/4 tablespoons

OpenStudy (calculusxy):

@KingGeorge @phi @e.mccormick

OpenStudy (calculusxy):

is it 5/6?

OpenStudy (kinggeorge):

So the original proportion is 2 cups of diced pumpkins to 1.25 tablespoons correct?

OpenStudy (calculusxy):

yes

OpenStudy (phi):

you need ⅔ of the original amount of sugar

OpenStudy (kinggeorge):

So that tells you that\[\frac{2}{1.25}=\frac{3}{x}\]where \(x\) is how many tablespoons you need for 3 cups of diced pumpkins.

OpenStudy (phi):

the original ratio was 3 / 1.25

OpenStudy (kinggeorge):

All you need to do from there, is solve for \(x\). Can you do that on your own?

OpenStudy (calculusxy):

can i just simply multiply 1.25/1 * 2/3= 2.5/3 = 5/6

OpenStudy (phi):

yes 2.5/3 is 5/6, which is correct.

OpenStudy (calculusxy):

okay can you please help me on my second problem please?

OpenStudy (calculusxy):

When Krish took the Saag Kumro to his friends, his friends start to savor the taste of the dish. Soon enough, Krish noticed that his friends had sent the news about how great the food was, to their friends. So 4 more people ask Krish to make the Saag Kumro with the same proportions of ingredients. Two want 1/2 of the dish, and the other two each want a whole bowl. How many more tablespoons of sugar does he need?

OpenStudy (kinggeorge):

I don't think 5/6 is correct...

OpenStudy (calculusxy):

i got the fact that i have to multiply 5/6*3 (since there is 3 cakes)

OpenStudy (kinggeorge):

Oh. I read the problem wrong...oops.

OpenStudy (phi):

King, the problem says Krish does not have 3 cups (which goes with the 1.25 tblsp of sugar)

OpenStudy (phi):

they never say how many bowls the recipe makes. Does he need 3 cups of pumpkin to make 1 bowl ? or just the 2 cups ?

OpenStudy (phi):

Or is a bowlful a cup ?

OpenStudy (calculusxy):

a bowlful is like 1. so 1/2+1/2=1 and 1+1=2, which makes it 1+2=3 all together.

OpenStudy (phi):

ok, so that means he has to make 3 times what the recipe calls for .

OpenStudy (calculusxy):

yes so i did 5/6*3=15/6

OpenStudy (phi):

or it could be 3 * 1.25 = 3.75 = 15/4 (using the original amount that the recipe called for). Unless they are clearer, I don't know what they want.

OpenStudy (phi):

I guess "with the same proportions of ingredients" might mean using the 5/6. I would go with your answer.

OpenStudy (calculusxy):

what about if krish uses the amount of the ingredients (like 5/6 of sugar and sticking with 2 cups of diced pumpkins)? how can i solve for it like that?

OpenStudy (phi):

you just did. 3*5/6 =15/6 = 5/2 tblsp of sugar

OpenStudy (calculusxy):

do i subtract? i did this problem by subtracting 15/6-5/6=10/6=1 2/3 <--- answer

OpenStudy (phi):

If you need to make 3 times the amount as before and he used 5/6 tblsp of sugar before, he will need 3*5/6 = 15/6 = 5/2 Are you asking a different question ?

OpenStudy (calculusxy):

i understood that. but what about the last statement "How many more tablespoons of sugar does he need?"

OpenStudy (phi):

oh, good point. yes, what you did looks good. He needs 10/6 = 1 ⅔ more than before

OpenStudy (calculusxy):

how can i write part b (second question) in a wording that says that he used 2 cups of diced pumpkins and 5/6 amount of sugar. basically, how can i make the second question depend on the first question?

OpenStudy (phi):

You could write Given 2 cups of pumpkin and 5/6 tablespoon of sugar makes one bowl of Saag Kumro. Then 3 bowls of Saag Kumro requires a total of 2 ½ tablespoons of sugar, or an additional 1 ⅔ tablespoons of sugar beyond the original 5/6 tablespoons.

OpenStudy (calculusxy):

and that would give me a result of 1 2/3 right?